bennys collection of nickles and dimes is worth $2.00. He has 22 total coins. How many of each type of coin does he have?


Answer by Abbeyneedshelp!(2)

Key- * <----Times Sign. (means multiply.)
]
Dimes/ 10*10=100 so 0.10 (cents) * 10 = $1.00
]
Nickles/ 0.05(cents) * 20 = $1.00
]
Benny Has 10 Dimes and 20 Nickles in his collection.